A Store-It-All storage facility requires users to enter a four digit code, using numbers 0-9, to get through the gate. You need to get a box of winter clothes from your storage container, but you forgot the code. 

  1. If the digits can be repeated, how many possible codes are there?

There are _____ different codes. 


1
Expert's answer
2021-02-24T12:09:18-0500

"N=10*10*10*10=10^4=10000."

There are _10000_ different codes.